OpenStudy (anonymous):

(2 cis (pi)) (8 cis (2 pi))

OpenStudy (anonymous):

1111111111111111113.314

OpenStudy (anonymous):

\[(2 cis (\pi)) \times (8 cis (2\pi))\]